OpenStudy (anonymous):

How are solutions, roots, and x-intercepts of a quadratic related? NEED HELP ASAP PLZ :D please explain I'm stuck

OpenStudy (anonymous):

HELP ME PLZZZZ

OpenStudy (anonymous):

Heeeeeelllllppppppp seriously mathematicians of the

OpenStudy (anonymous):

WORLD

OpenStudy (baru):

all of them mean the same thing

OpenStudy (anonymous):

@baru but i need an explanation :( they have different purposes but I need to know how they connect

OpenStudy (baru):

when we say: x intercept, its usually in a graph context that is y=\(ax^2+bx+c\) x intercepts are points where y=0

OpenStudy (baru):

solution is when we solve for 'x' \(ax^2+bx+c=0\) (find a value of x which fits that equation)

OpenStudy (baru):

so you see, both mean the same thing

OpenStudy (anonymous):

its still confusing but that helped :) but what about roots? how are they related to the solutions and intercepts. I thought the root was the solution to the discriminant.

OpenStudy (baru):

roots=solution=intercepts they are three different words that mean the same thing...

OpenStudy (baru):

discriminant helps you find the solution(or root or x intercept)

OpenStudy (anonymous):

ohhhhh kk then so the discriminant solution is the same as the x intercepts ?

OpenStudy (baru):

"discriminant" is a formula, using which you find solution/root/x intercept

OpenStudy (lochana):

all are same

OpenStudy (anonymous):

@baru ok then thanks :)

OpenStudy (anonymous):

@lochana i know but i needed an explanation :)

OpenStudy (lochana):

well, they all are same. roots are "a" and "b". where (x - a)(x - b) = 0 this is what we do with discriminant and all other things. we are trying to find "a" and "b" now go back to (x-a)(x-b) =0 in order to (x-a)(x -b) be zero, either (x-a) or (x - b) has to be zero. so x-a = 0 x - b = 0 that means x = a x = b is I draw this in a graph, "a" and "b" will be like this